OpenStudy (anonymous):

Find all roots of the equation 6y5 - 13y4 - 6y3 + 17y2 - 4 = 0

OpenStudy (anonymous):

help

OpenStudy (anonymous):

first you have to simplify to do anything else

OpenStudy (anonymous):

simplify and work it out take your time

OpenStudy (anonymous):

put y=1, 6*1^5-13*1^4-6*1^3+17*1^2-4=0 23-23=0 0=0 which is true. Hence y=1 is a root. y-1 is a factor of this eq. by synthetic division,  (y-1)(6y^4-7y^3-13y^2+4y+4)=0 \[6y ^{4}-7y ^{3}-13y ^{2}+4y+4=0\] y+1=0,gives y=-1 \[put y=2 \in6y ^{3}-13y ^{2}+4=0\] \[6*2^{3}-13*2^{2}+4=0,52-52=0,0=0\] Hence y=2 is a root.  \[6y ^{2}-y-2=0\] \[6y ^{2}-4y+3y-2=0\] 2y(3y-2)+1(3y-2)=0 \[\left( 3y-2 \right)\left( 2y+1 \right)=0,y=\frac{ 2 }{ 3 },\frac{ -1 }{ 2 }\] roots are 1,1,2,2/3,-1/2